# Query of CC: Unearthing Large Scale Domain-Specific Knowledge from Public Corpora

### v3

@@ Appendix A Model Training Details
In this study, we train all models with bf16 mixed precision and only data parallel. We employed the AMSP Chen et al. (2023), shard optimizer state across 8 cards to reduce communication overhead. Simultaneously, data parallelism was employed for training. To enhance throughput and reduce memory consumption, we introduced the Flash attention 2 Dao (2023) module. All models underwent training for 50,000 steps, with a global batch size of 4 million tokens per step, totaling 200 billion training tokens.
@@ Appendix B Evaluation
BIG-Bench hard Suzgun et al. (2023)
@@
Appendix C Collection implement details
In the implement section, we primarily discuss three components, the public corpora (C.1), retrieval engine (C.2), and the post-processing settings (C.3).
C.1 Public Corpora
With the development of large language models, public corpora have become increasingly rich, including the Pile, RedPajama, and Common Crawl. To ensure the diversity of data sources for retrieval, we constructed our retrieval corpus based on Common Crawl. Common Crawl is an open-source web crawler project containing all publicly available web pages from 2013 to the present. In theory, it encompasses a significant portion of the information present on the web. Due to hardware cost constraints, we utilized WARC format data from the years 2016 to 2023 to build our retrieval database. We performed extraction, filtering, and cleaning procedures on the data obtained from Common Crawl to ensure the quality of the retrieval dataset. The final retrieval database comprises several billion records, occupying a total of 50TB of disk space.
C.2 Retrieval Engine
@@
C.3 Post Processing
To enhance the quality of Knowledge Pile, inspired by Wenzek et al. (2020), we conducted data quality filtering and deduplication in the post-processing stage.

### v4

@@ Appendix A Model Training Details
In this study, we train all models with bf16 mixed precision and only data parallel. We employed the AMSP (Chen et al., 2023), shard optimizer state across 8 cards to reduce communication overhead. Simultaneously, data parallelism was employed for training. To enhance throughput and reduce memory consumption, we introduced the Flash attention 2 (Dao, 2024) module. All models underwent training for 50,000 steps, with a global batch size of 4 million tokens per step, totaling 200 billion training tokens.
@@ Appendix B Evaluation
BIG-Bench hard (Suzgun et al., 2023)
@@
Appendix C More Evaluation of Llama-QoC and Mistral-QoC
In order to compare the effects of $\mathsf{Retrieve}$-$\mathsf{Pile}$ on knowledge-related reasoning abilities, Figure 8 contrasts the improvement based on Llama2 and Mistral on MMLU, with a specific focus on performance across different subjects.
Figure 8: Left: the comparison of performance in AGIEval Benchmark between Mistral and Mistral-QoC. Right: The performance comparison of performance in MMLU.
From the perspective of Llama-QoC and Llama, following training with $\mathsf{Retrieve}$-$\mathsf{Pile}$, the average performance on MMLU for all categories increased by at least 10 points. This notable performance enhancement can be attributed in part to the initially lower performance of Llama, while the $\mathsf{Retrieve}$-$\mathsf{Pile}$ exhibits relatively rich knowledge content, playing a significant role in driving the improvement of Llama2’s performance.
In comparison to Llama, Mistral’s performance improvement is relatively moderate. The results indicate that in the STEM field, Mistral-QoC demonstrates a higher performance improvement compared to other fields, rising from 53.23 to 56.6. In contrast, performance improvements in other categories hover around one point. Two possible reasons account for this phenomenon. Firstly, Mistral exhibits strong comprehension abilities in disciplines such as humanities and social sciences, consistently scoring above 65, leaving relatively limited room for improvement. In contrast, the model’s score in the STEM field is only 53.23, suggesting a greater potential for performance enhancement. Secondly, $\mathsf{Retrieve}$-$\mathsf{Pile}$ has the largest and most extensive dataset in the STEM field, contributing to the more pronounced performance improvement of Mistral-QoC on MMLU-STEM.
Figure 8 compares the Performance of professional and academic exams covered in AGIEval. Even in the case of professional academic exams, $\mathsf{Retrieve}$-$\mathsf{Pile}$ similarly leads to substantial performance improvements. As observed in the figure, apart from a slight decline in the SAT-English expression test (from Mistral 68.93 to 65.53), other tasks exhibit noticeable improvements after continued training in $\mathsf{Retrieve}$-$\mathsf{Pile}$. Notably, SAT-Math rises from 31.36 to 45, which is a significant improvement for Mistral. Legal exams such as JEC-QA-CA and JEC-QA-KD also show significant enhancements, with the performance increasing from 18.2 and 15.3 to 26.1 and 15.3. For other law-related exam, LSAT is divided into three aspects: Law-Analytics (LSAT-AR), Law-Logic(LSAT-LR), and Law-Reading(LSAT-RC) improvements of 2.17, 8.82, and 7.07 point, respectively. For the LogiQA, Mistral-QoC achieves an accuracy of approximately 45.01%, nearly a 20% improvement compared to Mistral (39.32)
Appendix D Collection implement details
In the implement section, we primarily discuss three components, the public corpora (D.1), retrieval engine (D.2), and the post-processing settings (D.3).
D.1 Public Corpora
With the development of large language models, public corpora have become increasingly rich, including the Pile, RedPajama, and Common Crawl. Common Crawl is an open-source web crawler project containing all publicly available web pages from 2013 to the present. In theory, it encompasses a significant portion of the information present on the web. Due to hardware cost constraints, we utilized WARC format data from the years 2016 to 2023 to build our retrieval database. We performed extraction, filtering, and cleaning procedures on the data obtained from Common Crawl to ensure the quality of the retrieval dataset. The final retrieval database comprises several billion records, occupying a total of 50TB of disk space.
D.2 Retrieval Engine
@@
D.3 Post Processing
To enhance the quality of $\mathsf{Retrieve}$-$\mathsf{Pile}$, inspired by Wenzek et al. (2020), we conducted data quality filtering and deduplication in the post-processing stage.
@@ D.3 Post Processing
Appendix E Prompt of Query Expanding
==================  Prompt of Question Extension  ================== Suppose you are a question creator and your task is to create a new question based on the example question! Note that the new questions should have the same domain as the example questions, but be less frequent, of exactly the same length and difficulty as the example questions. You need to use your ingenuity to create a problem that is completely different from the given problem. Sample questions will be given after ###Given Question###. You need to write the newly created question after ###Created Question###. ###Given Question### [question]
==================  Prompt of Thought Generation  ================== Suppose you are a expert and your task is answer the given problem and tell me how to get the answer! You need to write the answer after the ###Answer### symbol. Please write the chain of thought after the ###COT### symbol. ###Given Question### [question]